Given the following information regarding an income producing property, determine the unlevered internal rate of return (IRR): expected holding period: five years; 1st year expected NOI: $89,100; 2nd year expected NOI: $91,773; 3rd year expected NOI: $94,526; 4th year expected NOI: $97,362; 5th year expected NOI: $100,283; debt service in each of the next five years: $58,444; current market value: $885,000; required equity investment: $221,250; net sale proceeds of property at end of year 5: $974,700; remaining mortgage balance at end of year 5: $331,026.
Answer #1

A B 1 2 Calculation of IRR Year Cash flow 0 -885000 5 1 89100 6 2 91773 7 3 94526 8 4 97362 9 5 1074983 10 IRR 12.21%

Excel formula:

2 Calculation of IRR 3 Year 40 5 1 62 73 84 95 Cash flow -885000 89100 91773 94526 97362 =100283+974700 =IRR(B4:B9) IRR